Translate to Algebraic symbols:
Krizia's age 5 years from now is she was x years old 2 years ago.

Answer by ramkikk66(644)

Translate to Algebraic symbols:

Krizia's age 5 years from now is she was x years old 2 years ago.
Ans:
If she was x years old 2 years ago,
current age = x + 2
Age 5 years from now = current age + 5 = (x + 2) + 5 = x + 7
:)